Problem 7.4: Use Eigen method and substitution method to find GS of X(e) = ({ -- ) x(0) + (%)** + (6)
Answer #1

The giren System is x0) = 2 47x0) + (%) et+ (0) + wine -E] Now, this system can be written as, .. e x = 2x + 4 X2 + t - 1 x= 15) (1-3)*(1+t-}) et - (-+) (1 + 2 + 2 . ) (++ ) - 2 et =(-) (++) - - (-1). (+5) - et + .: therefore, x-gest + C e 2t - 1 e